I do not understand how to fugure out these problems:
"The sum of an integer and its square is 30. Find the number." My choices are -3 or 10, 3 or -10, -5 or 6, and 5 or -6.
I am totally confused as to why there are negative numbers involved. Can someone please explain how to solve this problem step by step as I have others similar to this one that I am also confused on. Thank you!

Answer by stanbon(75887)

The sum of an integer and its square is 30.
-------
Let the integer be "x".
EQUATION:
x + x^2 = 30
x^2 + x - 30 = 0
Factor:
(x+6)(x-5) = 0
x = -6 or x = 5
=======================
Cheers,
Stan H.